OpenStudy (luigi0210):

Find the derivative:

OpenStudy (luigi0210):

\[y=4x \sqrt{x+\sqrt{x}}\]

OpenStudy (anonymous):

OK so do u know the basic Derivation formulas??

OpenStudy (bahrom7893):

Chain rule/product rule/power rule

OpenStudy (bahrom7893):

\[y=4x*(x+x^{1/2})^{1/2}\]

OpenStudy (anonymous):

I'll leave it to @bahrom7893 !!He knows wat he's doin!! -_<

OpenStudy (bahrom7893):

\[y' = 4*(x+x^{1/2})^{1/2}+4x*((x+x^{1/2})^{1/2})'\]

OpenStudy (bahrom7893):

So far so good?

OpenStudy (luigi0210):

Yup.

OpenStudy (bahrom7893):

Okay so now let's find the following: \[((x+x^{1/2})^{1/2})'=\frac{1}{2}(x+x^{\frac{1}{2}})^{-\frac{1}{2}} * (1+\frac{1}{2}x^{-\frac{1}{2}})\]

OpenStudy (bahrom7893):

Are you still following? I'll give you a moment to let it sink in.

OpenStudy (anonymous):

O_O!! Now dats wat I call a Derivative!!

OpenStudy (luigi0210):

Yup yup, all good here.

OpenStudy (bahrom7893):

Okay, so now just substitute whatever's in the parenthesis into the previous equation, and do some simplification.

OpenStudy (luigi0210):

So answer I got \[y=\frac{6x+5\sqrt{x}}{\sqrt{x+\sqrt{x}}}\]

OpenStudy (bahrom7893):

it's probably right.. sorry im too lazy to simplify.. @.Sam. can you check?

sam (.sam.):

His answer is right out of wolf, but not sure about you @bahrom7893 xD

OpenStudy (bahrom7893):

Well if it is right, then we're good to move on.

OpenStudy (luigi0210):

I didn't use Wolfram.

OpenStudy (luigi0210):

Anyways, thanks for your help guys!

sam (.sam.):

What I've got is \[y'=(4x)(\frac{1}{2})(x+x^{1/2})^{-1/2}(1+\frac{1}{2\sqrt{x}})+4\sqrt{x+\sqrt{x}}\]

sam (.sam.):

For x=2, y'=10.32, you can use this as a guide in case you have any mistakes